OpenStudy (anonymous):

Multiplying Polynomials (3xy - 1) (4xy + 2)

OpenStudy (anonymous):

\[(3xy-1)(4xy+2)\]\[3xy \times 4xy + 3xy \times 2 + -1 \times 4xy + -1 \times 2\]\[12x^{2}y^{2}+6xy-4xy-2\]\[12x^{2}y^{2}+2xy-2\] Other forms: \[2(3xy-1)(2xy+1)\]\[x(12xy^{2}+2y)-2\]OR\[y(12x^{2}y+2x)-2\]